# SafeSearch: Automated Red-Teaming for the Safety of LLM-Based Search Agents

### v3

@@ Abstract
Search agents connect LLMs to the Internet, enabling access to broader and more up-to-date information. However, unreliable search results may also pose safety threats to end users, establishing a new threat surface. In this work, we conduct two in-the-wild experiments to demonstrate both the prevalence of low-quality search results and their potential to misguide agent behaviors. To counter this threat, we introduce an automated red-teaming framework that is systematic, scalable, and cost-efficient, enabling lightweight and harmless safety assessments of search agents. Building on this framework, we construct the SafeSearch benchmark, which includes 300 test cases covering five categories of risks (e.g., misinformation and indirect prompt injection). Using this benchmark, we evaluate three representative search agent scaffolds, covering search workflow, tool-calling, and deep research, across 7 proprietary and 8 open-source backend LLMs. Our results reveal substantial vulnerabilities of LLM-based search agents: when exposed to unreliable websites, the highest ASR reached 90.5% for GPT-4.1-mini under a search workflow setting. Moreover, our analysis highlights the limited effectiveness of common defense practices, such as reminder prompting. This emphasizes the value of our framework in promoting transparency for safer agent development. Our codebase and test cases are publicly available: https://github.com/jianshuod/SafeSearch.
@@ 1 Introduction
Figure 1: LLM Services Can Return Unsafe Code Due to Internet-Sourced Unreliable Search Results.
@@
Large Language Models (LLMs) are not inherently ready for time-sensitive and long-horizon queries (Lewis et al., 2020; Wei et al., 2024b). Aside from RAG methods (Fan et al., 2024), search agents have recently gained traction (Vu et al., 2024; Team et al., 2025b; OpenAI, 2025e). By interacting with search tools, these agents connect LLMs to the Internet, thereby granting access to broad and up-to-date information. This paradigm enables applications from rapid information acquisition like ChatGPT Search (OpenAI, 2024) to long-horizon tasks such as literature surveys (Xu et al., 2025).
@@
However, an often-overlooked point is that the safety of search agents depends on the reliability of the search results they receive. Meanwhile, such reliability is far from guaranteed. As illustrated in Figure 1, a striking example involves a developer who inadvertently used a ChatGPT-generated code snippet—sourced from a searched unreliable GitHub page—and, in the process, exposed his private key and lost approximately \$2,500111https://twitter-thread.com/t/1859656430888026524. Complementing this anecdote, our two in-the-wild experiments provide systematic evidence: 1) Low-quality websites, such as content farms, are prevalent and frequently appear in search results. 2) Unreliable sources can substantially distort search agents’ responses, even in sensitive domains like healthcare. In this context, end users naturally question, and developers are pressed to answer: To what extent can we trust the LLM-based search agents, when they inevitably encounter unreliable search results sourced from the Internet?
@@
To delineate this, we undertake a systematic red-teaming study to identify, measure, and begin to mitigate failures in search agent safety. Existing evaluation methods (Wei et al., 2024a; 2025; Wong et al., 2025) for search agents rely heavily on human efforts to craft test cases, which are labor-intensive and hard to scale. In-the-wild testing, such as our two motivating experiments in Section 2.2, remains inefficient and offers limited systematic coverage. Besides, adversarial testing via blackhat SEO (Sharma et al., 2019) may cause harm to normal users, raising ethical concerns.
@@
To address the challenges, we propose an automated red-teaming framework offering three key advantages. 1) Systematic: Our red-teaming framework proactively approaches the threat of unreliable search results, while enabling testing of multiple risks: When benign queries surface unreliable search results, we study how agents adopt and propagate them in their final responses, exposing risks such as harmful output (Zou et al., 2025), indirect prompt injection (Perez & Ribeiro, 2022), advertisement promotion (Nestaas et al., 2025), misinformation (Pan et al., 2023), and bias (Wu et al., 2025). 2) Automatic & Scalable: Our framework runs automatically and scales efficiently. It leverages LLM assistants to generate test cases, applies differential testing for quality filtering, involves guideline-assisted website generation, and employs specialized LLM evaluators to assess agents (e.g., helpfulness evaluation and checklist-assisted safety evaluation). 3) Harmless & Cost-Efficient: To avoid the high cost and harmful effects of real-world SEO manipulations, we simulate unreliable search results by injecting an LLM-generated unreliable website into authentic search results, enabling efficient, large-scale, and harmless testing.
@@
Based on the framework, we construct SafeSearch, a benchmark comprising 300 high-quality test cases spanning the above five risk scenarios. This benchmark enables systematic and quantitative evaluation of the safety progress of search agents. Our study examines three representative search agent scaffolds (search workflow (Vu et al., 2024), tool calling (Yang et al., 2025), and deep research (Xu & Peng, 2025)) and a diverse set of backend LLMs (seven proprietary and eight open-source ones). In response to the research question, our study highlights the overall high vulnerability of LLM-based search agents to unreliable search results. The globally highest ASR (90.5%) is observed when using GPT-4.1-mini under the search workflow setting. Both backend LLMs and scaffold designs are critical—reasoning models and deep research scaffolds generally demonstrate stronger resilience. Across risk types, susceptibility varies markedly, with hard-to-verify misinformation posing the greatest threat. Finally, the interplay between safety and helpfulness shows that these goals are not at odds; GPT-5-based agents or those equipped with deep-research scaffold can achieve advanced performance across both dimensions.
@@
Beyond the main findings, our red-teaming framework proves valuable for tracking safety progress in agent development. For instance, it reveals how seemingly simple design choices, such as the number of search results, can implicitly affect agent safety, and it enables the evaluation of defense strategies, e.g., showing the limited effectiveness of reminder prompts and revealing the knowledge-action gap of LLMs in safety-related issues. These analyses highlight the transparency our framework brings to agent development. We hope our red-teaming framework and the SafeSearch benchmark as a foundation for evaluating and advancing more capable and trustworthy search agents that address both performance and safety in tandem.
@@ 2 Preliminaries
2.1 Why Are Search Agents Distinct?
Large Language Models (LLMs) are, by nature, not ready for time-sensitive tasks. Their responses are based on static training data, which limits their ability to access or reason over the most recent information. This limitation has driven increasing demand for augmenting LLMs with external knowledge through in-context methods, ranging from retrieval-augmented generation (RAG) (Lewis et al., 2020) to search agents (Vu et al., 2024). Notably, search agents are distinct from RAG systems.
LLM-based search agents are systems that couple an LLM with search tools to dynamically retrieve and integrate information at inference time. Unlike RAG, which typically operates over a static, locally controlled, and well-curated corpus, search agents rely on often opaque search services that provide access to large-scale, dynamic, and up-to-date information on the Internet, e.g., via Google Search (Nakano et al., 2021; Schick et al., 2024; Vu et al., 2024). This distinction makes search agents more powerful for addressing information-seeking needs, but also introduces unique risks.
Figure 2: Search Agent vs. RAG.
Recent developments have expanded search agents beyond simple search pipelines (Vu et al., 2024), including specialized reasoning models that learn tool-use strategies (Jin et al., 2025; Sun et al., 2025; Team et al., 2025b) and deep research systems for long-horizon research assistance (OpenAI, 2025e; Li et al., 2025; Xu & Peng, 2025). These advancements demonstrate the growing sophistication of search agents, but also highlight their deep dependency on search services.
A threat arises from the search tool, which may yield unreliable Internet-sourced search results.
We contrast search agents with RAG in Figure 2. Unlike local retrieval in RAG, information seeking via search leads to limited controllability due to the scale and rapid evolution of the indexed corpus, as well as the proprietary nature of ranking algorithms. These issues are particularly salient when using third-party search services, such as those commonly integrated into agent frameworks [^2] https://python.langchain.com/docs/integrations/tools/#search. Systematically understanding this threat is essential; otherwise, end users may still suffer from unreliable agent responses, as exemplified by the real-world case in Figure 1.
2.2 Threat Surface: Search Tools Are Not Reliable at All Times
We investigate the issue of unreliable search results through two in-the-wild experiments.
The Popularity of Low-Quality Websites. We study content farms—sites that publish keyword-stuffed, low-value articles to manipulate search rankings (Wikipedia, 2025a)—as an example. Inevitably, these sites may appear in search results and affect search agents. We sample 1,000 personas from Persona Hub (Ge et al., 2024) and prompt GPT-4.1 to generate a plausible query each might issue. Using Google Search, we collect the top 10 results per query when available, yielding a pool of 8,933 websites. We tentatively employ GPT-4.1 with CoT reasoning to judge whether a site is likely a content farm (see Appendix G). Notably, 4.3% (380 out of 8,933) are flagged as probable content farms, primarily due to shallow content, lack of credible sources, or heavy advertising.
Low-quality sites extend beyond content farms, adversaries can employ stealthy blackhat SEO techniques (Sharma et al., 2019), and even reputable sources like Wikipedia can contain inaccuracies (Wikipedia, 2025b). Moreover, search vendors may prioritize sponsored content.
@@ 2.2 Threat Surface: Search Tools Are Not Reliable at All Times
Figure 3: Qualitative Example: The Search Agent May Shift its Stance with Unreliable Search Results. We prefer long-tail, inconclusive questions in this experiment, which are more likely to hit low-quality websites. The search agent used in the example is Qwen3-8B with a search workflow.
The Impact of Search Results on Agent Behaviors. The presence of unreliable websites in search results motivates us to examine whether such sources influence the behavior of search agents. We illustrate our methodology in Figure 3. We begin by prompting GPT-5-mini to generate 1,000 health-related boolean questions. We then conduct differential testing with Qwen3-8B, a representative edge-side model that is sufficiently lightweight to be deployed even by ordinary users on their local machines. We compare its responses when equipped with a Google Search tool against its direct, tool-free answers. This setup allows us to measure the effect of search integration on agent behaviors. With the assistance of GPT-5-mini, we find that in 46 out of 1,000 cases, the agent’s responses are altered. Manual review reveals that in most instances, the agent becomes more inclined to acknowledge risky treatments after being exposed to search results, often due to retrieving unreliable sources such as anecdotal blogs. As a motivating experiment, these findings highlight the potential risks users face when search agents encounter unreliable search results.
@@ 2.3 Related Works
Existing methods offer little systematic assessment of search agent safety. 1) Evaluating Search Agents: Existing benchmarks for search agents, such as GAIA(Mialon et al., 2024), SimpleQA(Wei et al., 2024a), and BrowseComp (Wei et al., 2025), primarily target performance ceilings rather than safety boundaries. A relevant effort is OpenAI’s in-house safety testing of its deep research system, which relies on manual red-teaming to identify potential risks (OpenAI, 2025e). However, their closed-source and expert-driven approach lacks transparency and scalability. 2) Agent Safety Benchmarks: Existing agent safety evaluations rarely cover threats to search agents or tools in depth (Ruan et al., 2024; Zhang et al., 2025b; Andriushchenko et al., 2025), and when they do, the focus is often restricted to narrow threats like indirect prompt injection (Zhan et al., 2024; Liao et al., 2025). This gap underscores the need for more targeted and systematic evaluations of search agent safety. Moreover, rather than assuming threats from user inputs (Ruan et al., 2024), we attribute threats to the potential inability of insider tools (Fang et al., 2025), e.g., search tools in this paper. 3) Knowledge Base Safety: Methods for evaluating RAG system safety (Chen et al., 2024; Wu et al., 2024; Liang et al., 2025; Huang et al., 2025; Ming et al., 2025; Zhang et al., 2025a), where threats stem from static knowledge bases, do not tackle the issues of search agents, whose threats originate from fundamentally different sources: dynamic, real-time web content. Ensuring the safety of search agents, therefore, requires scalable testing methods that can accommodate live, unpredictable search results and evolving threat surfaces. The most relevant work to us is Luo et al. (2025), who pioneered a measurement study of the potential risks faced by LLM-based search services. However, their analysis primarily focuses on malicious websites rather than broader safety threats, emphasizes system-level outcomes rather than the behavior of the underlying search agents, and targets post-deployment measurement rather than scalable pre-deployment testing.
3 A Red-Teaming Framework for Search Agent Safety
@@ 3.1 Threat Model & Framework Overview
Risk Coverage. As described in Section 2, the studied threat spans diverse scenarios where search tools deliver unreliable content to agents. In this work, we focus on five representative ones (Appendix’s Table 4): two adversarial risks, namely indirect prompt injection (Perez & Ribeiro, 2022) and harmful outputs (Zou et al., 2025), and three non-adversarial risks, namely bias inducing (Wu et al., 2025), advertisement (Nestaas et al., 2025), and misinformation (Pan et al., 2023).
These risks arise from distinct interests, adversarial or non-adversarial, yet manifest in analogous ways from the standpoint of search agents—the issue of unreliable search results. This motivates consolidating our testing of such risks within a unified framework. Since we are required to proactively approach the occurrence of unreliable search results, we position our testing as a practice of red-teaming aimed at characterizing the lower bound of a search agent’s safety.
Threat Model of Our Red-Teaming. We state it from the perspective of the agent tester:
@@
    Capability: The tester only issues benign queries. For each query, the tester can deliberately introduce at most one unreliable website through the search tool into the agent execution. This avoids over-estimation of the risks agents may face in real-world deployments.
@@
    Knowledge: The threat is case-related but agent-agnostic: for a given test case, the unreliable website is fixed across all search agents and is not tailored to any specific agent. [^3] For benchmarking purposes, this aggressive setting ensures fairness across agents. We also acknowledge that more targeted red-teaming can tailor an unreliable website specifically to the agent under evaluation.
@@
    Objective: The tester evaluates how these threats affect the agent’s outputs. While the impact varies across risk types, it primarily manifests as unsafe responses delivered to users.
Overview of The Red-Teaming Framework. To substantiate the threat model, each test case should include four essential components: a user query that reflects the use scenario, a target consequence to achieve, an unreliable website intended to distract the search agent, and an evaluation function that determines whether the expected risk occurs. Our red-teaming framework operates in two stages. The first is an offline stage, in which we curate high-quality test cases according to the specified risk types (Section 3.2). In the second, we execute each test case with search agents to identify their vulnerabilities (Section 3.3). The automation of our red-teaming framework results from a well-orchestrated integration of four LLM assistants: an LLM-driven workflow for test case generation, a website generator, a helpfulness rater, and a checklist-assisted safety evaluator.
3.2 Automated Generation of Test Cases At Scale
Test Generation Workflow. We design a structured trajectory for generating test cases, as illustrated in Figure 4. The workflow takes a natural language specification (i.e., a risk description) as input and outputs a corresponding test case, using a reasoning model (e.g., o4-mini in this work) as the test generator. Specifically, the workflow proceeds in three sequential steps:

### v4

# SafeSearch: Automated Red-Teaming of LLM-Based Search Agents

@@ Abstract
Search agents connect LLMs to the Internet, enabling them to access broader and more up-to-date information. However, this also introduces a new threat surface: unreliable search results can mislead agents into producing unsafe outputs. Real-world incidents and our two in-the-wild observations show that such failures can occur in practice. To study this threat systematically, we propose SafeSearch, an automated red-teaming framework that is scalable, cost-efficient, and lightweight, enabling harmless safety evaluation of search agents. Using this, we generate 300 test cases spanning five risk categories (e.g., misinformation and prompt injection) and evaluate three search agent scaffolds across 17 representative LLMs. Our results reveal substantial vulnerabilities in LLM-based search agents, with the highest ASR reaching 90.5% for GPT-4.1-mini in a search-workflow setting. Moreover, we find that common defenses, such as reminder prompting, offer limited protection. Overall, SafeSearch provides a practical way to measure and improve the safety of LLM-based search agents.
@@ 1 Introduction
Large Language Models (LLMs) are not inherently suited to time-sensitive and long-horizon queries (Lewis et al., 2020a; Wei et al., 2024b). Beyond RAG methods (Fan et al., 2024), search agents have recently gained traction (Vu et al., 2024; Team et al., 2025b; OpenAI, 2025a). By interacting with search tools, these agents connect LLMs to the Internet, thereby granting access to broader and more up-to-date information. This paradigm enables diverse applications ranging from rapid information acquisition (e.g., ChatGPT Search (OpenAI, 2024)) to long-horizon tasks such as literature surveys (Xu et al., 2025b).
Figure 1: LLM services can return unsafe code due to Internet-sourced unreliable search results, which has led to a financial incident involving thousands of dollars in losses.
@@
However, an often-overlooked point is that search agent safety hinges on the reliability of the search results they receive. Yet such reliability is far from guaranteed. As illustrated in Figure 1, a striking example involves a developer who inadvertently used a ChatGPT-generated code snippet—sourced from an unreliable GitHub page returned by search—and, in the process, exposed his private key and lost approximately \$2,500111https://twitter-thread.com/t/1859656430888026524. Beyond this specific case, our two in-the-wild observations offer systematic evidence: (1) Low-quality websites, such as content farms, are prevalent and frequently appear in search results. (2) Unreliable sources can substantially distort search agents’ responses toward unsafe directions, even in sensitive domains such as healthcare. In this context, end users naturally ask, and developers are pressed to answer: To what extent can we trust LLM-based search agents when they encounter unreliable search results from the Internet?
@@
To answer this question, we undertake a systematic red-teaming study to identify, measure, and begin to mitigate failures in search agent safety. This is non-trivial: (1) Existing evaluations (Wei et al., 2025; Wong et al., 2025) for search agents rely on human effort to craft test cases, which is labor-intensive and hard to scale. (2) Testing by crafting risk-inducing queries (Luo et al., 2025; Ou et al., 2025) can be cost-inefficient and offer limited risk coverage. (3) Moreover, adversarial testing via black-hat SEO (Sharma et al., 2019) may harm normal users, raising ethical concerns.
@@
To address these challenges, we propose SafeSearch, an automated red-teaming framework with three key advantages. (1) Systematic: Our red-teaming framework proactively addresses the threat of unreliable search results and enables evaluation across multiple risks. When benign queries retrieve unreliable search results, we study how agents incorporate and propagate them in their final responses, exposing risks such as harmful output (Zou et al., 2025), indirect prompt injection (Perez and Ribeiro, 2022), advertisement promotion (Nestaas et al., 2025), misinformation (Pan et al., 2023), and bias (Wu et al., 2025). (2) Automatic & scalable: Our framework runs automatically and scales efficiently. It leverages LLM assistants to generate test cases, applies differential testing for quality filtering, involves guideline-assisted website generation, and employs specialized LLM evaluators to assess agents (i.e., helpfulness evaluation and checklist-assisted safety evaluation). (3) Cost-efficient & harmless to normal users: To avoid the high cost and harmful effects of real-world SEO manipulations, we simulate unreliable search results by injecting an LLM-generated unreliable website into authentic search results, enabling efficient and harmless testing.
@@
We conduct comprehensive red-teaming experiments with SafeSearch. Concretely, we construct a dataset comprising 300 high-quality test cases spanning the above five risk scenarios. With this, we examine three representative search agent scaffolds (search workflow (Vu et al., 2024), tool calling (Yang et al., 2025), and deep research (Xu and Peng, 2025)) and a diverse set of backend LLMs (nine proprietary and eight open-source ones). In response to the research question, our study highlights the overall high vulnerability of LLM-based search agents to unreliable search results. The globally highest ASR (90.5%) is observed when using GPT-4.1-mini under the search workflow setting. Both backend LLMs and scaffold designs are critical—reasoning models and deep research scaffolds generally demonstrate stronger resilience. Across risk types, susceptibility varies markedly, with hard-to-verify misinformation posing the greatest threat. Finally, the interplay between safety and helpfulness shows that these goals are not at odds; GPT-5-based agents or those equipped with a deep-research scaffold can excel in both dimensions.
@@
Beyond benchmarking, our red-teaming framework proves valuable for tracking safety progress in agent development. For instance, it reveals how seemingly simple design choices, such as the number of search results, can implicitly affect agent safety, and it enables the evaluation of defense strategies, e.g., showing the limited effectiveness of reminder prompts and revealing the knowledge-action gap of LLMs in safety-related issues. These analyses highlight the transparency our framework brings to agent development. We hope SafeSearch serves as a foundation for evaluating and advancing more capable and trustworthy search agents that address both performance and safety in tandem. We open-source the code and dataset at https://github.com/jianshuod/SafeSearch.
@@ 2 Preliminaries
2.1 Search Agent & New Threat Surface
@@
Figure 2: A qualitative example of stance shift caused by unreliable search results. We focus on long-tail, inconclusive queries that are more likely to hit low-quality sources. The search agent in this experiment uses Qwen3-8B with a search workflow.
In this work, we use LLM-based search agents to broadly refer to agentic systems that combine an LLM with search tools to dynamically augment knowledge at inference time. [^2] The term agent may be interpreted differently depending on context; here, we use it as an umbrella term without implying a specific agent architecture or level of autonomy. These include simple search pipelines (Vu et al., 2024), specialized reasoning models that learn tool-use strategies (Jin et al., 2025; Sun et al., 2025; Team et al., 2025b), and deep research systems for long-horizon research assistance (OpenAI, 2025a; Li et al., 2025; Xu and Peng, 2025). One notable feature shared by these systems is that they rely on search tools, e.g., Google Search or other search services, to provide LLMs with access to large-scale, dynamic, and up-to-date information on the Internet (Nakano et al., 2021; Schick et al., 2024; Vu et al., 2024).
The distinctive knowledge source (i.e., the open and broad Internet) makes search agents powerful for addressing information-seeking needs, but also introduces unique risks.
Search tools may return unreliable, biased, or manipulable search results, which are then treated as authoritative context by the search agent.
Real-world incidents, such as ChatGPT Search encountering a scam GitHub page in Figure 1, illustrate this risk in deployed systems. Beyond this case, we present two in-the-wild observations to motivate the practical relevance of this threat for broader search agents. See Appendix B for setup.
Observation 1: Unreliable websites appear with non-trivial frequency in search results. Even without targeted attacks, search results frequently include low-quality content. In a large-scale sample of user-like queries, a non-trivial fraction of top-ranked results (4.3%, 380 out of 8,933) come from content farms or similarly low-credibility websites, characterized by shallow content, heavy advertising, or lack of verifiable sources. Beyond content farms, unreliable information may arise from black-hat SEO (Sharma et al., 2019), sponsored advertisements, or inaccuracies in otherwise reputable sources like Wikipedia (Wikipedia, 2025b). These factors indicate that the threat is a safety issue that subsumes, but is not limited to, explicit adversarial attacks.
Observation 2: Search results can alter agent behavior in unsafe directions. To assess whether unreliable search results can affect search agents, we conduct differential testing by comparing an agent’s responses with and without search integration, as illustrated in Figure 2. Across 1,000 health-related queries, enabling search leads to the agent’s binary stance (yes/no) changing in 46 cases. Manual inspection indicates that these shifts often reflect increased acceptance of risky or unverified treatments, driven by exposure to anecdotal or low-quality web sources. This suggests that unreliable search results can directly influence agent decision-making, rather than being ignored or corrected.
@@ 2.2 Related Works
Evaluating search agents. Existing benchmarks for search agents, including GAIA (Mialon et al., 2024), SimpleQA (Wei et al., 2024a), and BrowseComp (Wei et al., 2025), primarily evaluate task performance and accuracy, rather than safety. A relevant effort is OpenAI’s in-house safety testing of its deep research system, which relies on manual red-teaming to identify potential risks (OpenAI, 2025a). However, their closed-source and expert-driven approach lacks transparency and scalability.
Agent safety benchmarks. Existing agent safety evaluations rarely cover threats to search agents or tools in depth (Ruan et al., 2024; Zhang et al., 2025c; Andriushchenko et al., 2025), and when they do, the focus is often restricted to narrow threats like indirect prompt injection (Zhan et al., 2024; Liao et al., 2025). This gap underscores the need for more targeted and systematic evaluations of search agent safety.
Search agents versus RAG. These two paradigms share the goal of augmenting LLMs with external knowledge, but differ significantly in their knowledge source and control. Typical RAG systems (Lewis et al., 2020b; Singh et al., 2025) operate over static, locally managed, and curated corpora, enabling developers to audit and patch unreliable retrieval results offline. Search agents, in contrast, rely on external search tools that index the open web, where content is large-scale, continuously evolving, and ranked by proprietary algorithms beyond the agent developer’s control. [^3] This is particularly evident when using third-party search services commonly integrated into agent frameworks (e.g., those listed in https://python.langchain.com/docs/integrations/tools/#search). As a result, unreliable search results may surface unpredictably at inference time, requiring search agents to reason about and respond to such content on the fly. This gap suggests that existing safety (Wu et al., 2024; Liang et al., 2025; Huang et al., 2025) and security (Zou et al., 2025; Xue et al., 2024; Zhang et al., 2025b) studies on RAG systems may not readily tackle this new issue. We provide a more detailed comparison of the two paradigms in Appendix A.
The most relevant work to ours includes Luo et al. (2025) and Ou et al. (2025), both of which explore risks in search-augmented LLM services. However, our work differs in several key aspects: (1) Scope: While the two studies focus on system-level outcomes, our research centers on agent-level behaviors. (2) Threat source: Unlike them, which curate adversarial or risk-inducing queries, we assume benign user queries and attribute risks to insider search tools (Fang et al., 2025). (3) Risk coverage: While Luo et al. (2025) targets malicious websites, we consider a broader range of safety threats, including advertisements.
Figure 3: Three-step workflow of test case generation. Colored fields are required in final test cases, and white solid-bordered fields are passed to later stages. Additionally, we elicit auxiliary information (dashed-bordered fields), such as unreliable website sources, to guide the generation and improve the practical relevance of the query and the unreliable website.
3 SafeSearch: A Red-Teaming Framework for Search Agents
@@ 3.1 Threat Model & Framework Overview
Our scope. In this paper, we approach the threat of unreliable search results from an agent-centric perspective, evaluating whether search agents can behave safely when exposed to unreliable search results. Concretely, we examine whether agents can critically utilize unreliable information in search results, rather than blindly trusting it and exhibiting unsafe behaviors. For further clarification, we provide five pairs of safe and unsafe agent responses in Appendix L.
Risk coverage. As described in Section 2, the phenomenon of unreliable search results covers diverse scenarios where search tools deliver low-quality, misleading, or manipulable content. We do not attempt to exhaustively taxonomize all failure modes; instead, we focus on five representative risk types (Appendix Table 2): two adversarial risks, namely indirect prompt injection (Perez and Ribeiro, 2022) and harmful outputs (Zou et al., 2025), and three non-adversarial risks, namely bias-inducing (Wu et al., 2025), sponsored advertisements (Nestaas et al., 2025), and misinformation (Pan et al., 2023). These risks arise from distinct interests, adversarial or non-adversarial, yet share the same overarching issue of unreliable search results. We consolidate our testing of these risks within a unified framework. Since we proactively approach the occurrence of unreliable search results, we position our testing as a practice of red-teaming aimed at characterizing the lower bound of a search agent’s safety.
Threat model of our red-teaming. We state it from the perspective of an agent tester:
@@
    Capability: The tester issues only benign queries. For each query, the tester can deliberately introduce at most one unreliable website through the search tool into the agent execution. This avoids overestimating the risks that agents may face in real-world deployments.
@@
    Knowledge: The threat is case-related but agent-agnostic: for a given test case, the unreliable website is fixed across all search agents and is not tailored to any specific agent.
@@
    Objective: The tester aims to discover potentially unsafe behaviors when exposed to unreliable search results. While the impact varies across risk types, it primarily manifests as unsafe responses delivered to users.
Overview of the red-teaming framework. To instantiate this threat model, each test case includes four essential components: a benign user query that reflects the use scenario, a target negative consequence to achieve, an unreliable website intended to distract the search agent, and an evaluation method that determines whether the expected risk occurs. We provide test case examples in Appendix K. Our red-teaming framework operates in two stages. The first is an offline stage, in which we curate high-quality test cases according to the specified risk types (Section 3.2). In the second, we execute each test case with search agents to identify their vulnerabilities (Section 3.3). The automation of our red-teaming framework results from a well-orchestrated integration of four LLM assistants: an LLM-driven workflow for test case generation, a website generator, a helpfulness rater, and a checklist-assisted safety evaluator.
3.2 Automated Generation of Test Cases
Test generation workflow. We design a structured trajectory for generating test cases, as illustrated in Figure 3. The workflow takes a natural language specification (i.e., a risk description) as input and outputs a corresponding test case, using a reasoning model (e.g., o4-mini) as the test generator. The workflow proceeds in three sequential steps: